OpenStudy (anonymous):

use formulas to find the lateral area and surface area of the given prism. The base of the prism is a rectangle with sides 10 m and 5 m in length. Round the answers to the nearest quare meter.

OpenStudy (jhannybean):

Lateral Surface Area \[\large L = 2(10 \cdot 5) + 2(5\cdot 5)\]\[\large L = 2(50) + 2(25)\ m^2\] Surface area  \[\large SA= 4 ( 10 \cdot 5) +2(5\cdot 5)\]\[\large SA= 4(50)+2(25) \ m^2\]